A 0.800 mm -diameter ball bearing has 2.20×109 excess electrons. Part A What is the ball bearing's potential?
Answer #1

charge on bearing

Q = ne = 2.2*10^9*1.6*10^-19 = 3.52*10^-10 C

potential

V = k Q / r

V = 9*10^9*3.52*10^-10 / (0.4*10^-3)

V = 7920 V

=======

do comment in case any doubt, will reply for sure . Goodluck